# Modeling reflection spectra of super-Eddington X-ray sources

### v1

@@ 2 Model description
RELLINE_SD $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$ $\times$
REFLUX $\times$ $\times$ 15 $\times$ $\times$
REFLUXCp $\times$ $\times$ 15 $\times$ $\times$
REFLUXD $\times$ $\times$ 300 $\times$ $\times$
RELLINELP_SD $\times$ $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$ $\times$
REFLUXLP $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$
REFLUXLPCp $\times$ $\times$ $\times$ $\times$ 15 $\times$ $\times$
REFLUXLPD $\times$ $\times$ $\times$ $\times$ 300 $\times$ $\times$
Table 1: Parameter availability for different sub-models. Here, ✓means the parameter is available in the given sub-model and $\times$ means the parameter is not available.
3 Spectral fitting of a ULX pulsar
In this section, we discuss the results of our fitting we obtained for the Galactic ULX pulsar Swift J0243.6+6124 (Cenko et al., 2017; Kennea et al., 2017). The source went into a bright outburst in 2017, with its peak luminosity $>10^{39}$ erg s-1, far exceeding the Eddington limit for a neutron star (Wilson-Hodge et al., 2018).
We analyze the quasi-simultaneous NICER (ObsID 1050390113, 1 November 2017) and NuSTAR (ObsID 90302319004, 30 October 2017) observations that have been fitted with the standard thin-disk reflection model in Jaisawal et al. (2019) (see their Fig. 7). We follow the procedure in Jaisawal et al. (2019) to process the NuSTAR (FPMA and FPMB) data. The NICER data are first processed with nicerl2 (CALDB v20240206 and geomagnetic data updated to March 2024). The source and background spectral files are then extracted using nicel3-spect with the default background model. Spectra are grouped to ensure a minimal counts of 100 for each bin before fitting with XSPEC v12.13.0c (Arnaud, 1996). Best-fit parameter values and uncertainties (90% confidence level unless otherwise stated) are obtained using $\chi^{2}$ statistics.
As in Jaisawal et al. (2019), we first fit the NICER (1–10 keV) and NuSTAR (3–79 keV) data with the model: constant * tbabs * (bbodyrad + cutoffpl + relxill/relxilllp). In this model, the relxill and relxilllp (García et al., 2014) component represent relativistic reflection from a razor-thin accretion disk, assuming fixed emissivity ($q=3$) and the corona as a point source above the black hole at a certain height ($h$), respectively. We fix the spin parameter at $a_{*}=0$ and the outer radius of the accretion disk at 1000 $M$. The best-fit parameters and the corresponding uncertainties are shown in Tab. 2 for fixed emissivity and in Tab. 3 with lamp-post corona geometry.
At the time of this observation, the mass accretion rate of the source was at super-Eddington level. The absorption corrected X-ray flux in the 0.01-100 keV band is around 10 times of the Eddington limit. Therefore, a thick disk, instead of thin one, is expected. We replace the relxill/relxilllp component with our REFLUX/REFLUXLP model, with the mass accretion rate fixed at 10 times of the Eddington accretion rate. As shown in Tabs. 2 and 3, the models provide a good fit to the data, with the $\chi^{2}$ close to that of relxill/relxilllp. The best-fit model components and residuals are shown in Fig. 4.
[…]
Tbabs nH $10^{22}$ cm-2 $0.683_{-0.011}^{+0.009}$ $0.664_{-0.006}^{+0.01}$
Cutoffpl $\Gamma$ $1.423_{-0.009}^{+0.009}$ $1.434_{-0.009}^{+0.014}$
Ecut keV $24.3_{-0.5}^{+0.6}$ $24.8_{-0.4}^{+0.3}$
norm $12.4_{-0.5}^{+0.5}$ $12.40_{-0.29}^{+0.5}$
Bbodyrad kT keV $1.213_{-0.018}^{+0.019}$ $1.234_{-0.013}^{+0.009}$
norm $955_{-33}^{+21}$ $988_{-27}^{+25}$
Reflection Incl deg $15.1_{-2.4}^{+4}$ $25.5_{-1.7}^{+3}$
$R_{\rm in}$ $M$ $37_{-7}^{+9}$ $34_{-7}^{+8}$
$\log(\xi)$ erg cm s-1 $3.54_{-0.07}^{+0.05}$ $3.29_{-0.12}^{+0.03}$
$A_{\rm Fe}$ solar $4.2_{-1.2}^{+1.4}$ $2.39_{-0.27}^{+0.5}$
norm $0.050_{-0.005}^{+0.007}$ $0.0255_{-0.0017}^{+0.001}$
$C_{\rm FPMB}$ $0.9880_{-0.0014}^{+0.0014}$ $0.9880_{-0.0014}^{+0.0014}$
$C_{\rm XTI}$ $1.242_{-0.012}^{+0.012}$ $1.240_{-0.012}^{+0.01}$
$\chi^{2}/\nu$ 2855.14/2624 2857.12/2624
Table 2: Best-fit parameters for the model with relxill and the model with REFLUX. The reported uncertainties correspond to the 90% interval limits for one relevant parameter ($\Delta\chi^{2}=2.71$).
[…]
Tbabs nH $10^{22}$ cm-2 $0.667_{-0.009}^{+0.009}$ $0.671_{-0.008}^{+0.007}$
Cutoffpl $\Gamma$ $1.434_{-0.009}^{+0.01}$ $1.438_{-0.01}^{+0.008}$
Ecut keV $24.74_{-0.4}^{+0.26}$ $24.7_{-0.3}^{+0.3}$
norm $12.4_{-0.3}^{+0.4}$ $12.6_{-0.3}^{+0.3}$
Bbodyrad kT keV $1.23_{-0.01}^{+0.01}$ $1.235_{-0.006}^{+0.005}$
norm $1000_{-30}^{+40}$ $970_{-25}^{+30}$
Reflection $h$ $M$ $2.20_{-P}^{+0.07}$ $>65$
Incl deg $25.7_{-1.3}^{+2.7}$ $26.5_{-2.4}^{+1.2}$
$R_{\rm in}$ $M$ $39_{-8}^{+9}$ $<14$
$\log(\xi)$ erg cm s-1 $3.31_{-0.05}^{+0.04}$ $3.284_{-0.07}^{+0.029}$
$A_{\rm Fe}$ solar $2.4_{-0.3}^{+0.3}$ $2.6_{-0.3}^{+0.4}$
norm $3.58_{-0.7}^{+0.24}$ $1.37_{-0.06}^{+0.10}\times 10^{-7}$
$C_{\rm FPMB}$ $0.9880_{-0.0014}^{+0.0014}$ $0.9880_{-0.0014}^{+0.0014}$
$C_{\rm XTI}$ $1.223_{-0.013}^{+0.012}$ $1.238_{-0.01}^{+0.011}$
$\chi^{2}/\nu$ 2816.49/2623 2826.35/2623
Table 3: Best-fit parameters for the model with relxilllp and the model with REFLUXLP. The reported uncertainties correspond to the 90% interval limits for one relevant parameter ($\Delta\chi^{2}=2.71$).
Figure 4: Best-fit model components and residuals for the lamppost geometry (left) and the power-law emissivity profile (right) with the REFLUX model. The black and red colors represent data of FPMA and FPMB, respectively. NICER data are plotted in cyan. Data are rebinned for visual clarity.
@@ 4 Discussions
The quality of the fits with the model with relxill and the model with REFLUX is similar. We fit the same data and the two models have the same number of free parameters, so we can directly compare their $\chi^{2}$ and the difference is marginal ($\Delta\chi^{2}=1.98$ for fixed emissivity and $9.86$ for lamppost). The $\chi^{2}$ of the best-fit with relxill is slightly lower – a possible reason for this could be the smaller parameter grid of REFLUX than relxill in the FITS file. However, we know that the source was accreting around 10 times of its Eddington limit, so the fit with REFLUX appears to be more appropriate for the spectra analyzed. In Tab. 2, the two models provide consistent estimates of the parameters with the exception of the inclination angle of the disk and the iron abundance. The measurements of the inclination angle are $i=15.1_{-2.4}^{+4}$ deg (relxill) and $25.5_{-1.7}^{+3}$ (REFLUX). However, the estimates of disk inclination angles inferred from the analysis of reflection features are normally not very accurate and both models suggest a low value of $i$. Estimates of the iron abundance using X-ray reflection spectroscopy are even more problematic. It is currently unclear why reflection measurements often suggest super-solar iron abundances (but there are also sources for which reflection measurements suggest iron abundances significantly lower than the solar one); see, for instance, Bambi et al. (2021) for a short review on the problem. It may be related to a deficiency of current reflection models (e.g., underestimate of the disk electron density (Jiang et al., 2019a, b; Liu et al., 2023b), presence of two or more coronae (Fürst et al., 2015), etc.) or a true physical phenomenon, like radiative levitation of metal ions in the inner part of the accretion disk (Reynolds et al., 2012). The estimate of the iron abundance with REFLUX would go to the right direction, decreasing the value of $A_{\rm Fe}$ with respect to that inferred with relxill, but we would need to analyze more sources and more observations to see if this is a feature of our model and not an accidental result for the spectra analyzed in the previous section. In Tab. 3, we see that for a lamppost geometry, the inclination angle and the iron abundance for both the models closely match the result of REFLUX in the analysis with fixed emissivity. However, we see a significant difference in the corona height and the inner edge of the disk. With relxill, we see a very low corona height (the lowest in the model) and a larger value for the inner edge. With REFLUX, the case is the opposite, we obtain a higher corona height $h>65M$ and inner edge closer to the central object $R_{in}<14M$. In the case of REFLUX, we see a small norm possibly because the whole disk may not be illuminated for many different $h$ and $R_{in}$ configurations making the flux from the outer part of the disk negligible. We caution the readers against interpreting the high value of $h>65M$ too rigidly. As in our analysis, we assume a lamppost geometry, which may not be applicable to an Ultra-Luminous Pulsar (ULP) like Swift J0243.6+6124. The system might instead contain a more extended primary illuminating source, such as an accretion column (e.g., Pinto & Walton (2023)). Nonetheless, we observed changes in inferred geometries when a slim disk model was employed. This underscores the critical importance of considering disk geometry when addressing systematic uncertainties. Such uncertainties significantly affect the inferred disk inner radius and, consequently, the magnetic field of ULPs (e.g., Jaisawal et al. (2019)). We will further investigate Swift J0243.6+6124 in this perspective in the future.
A more detailed analysis of the 2017 NICER and NuSTAR data of Swift J0243.6+6124 is beyond the purpose of the present work, which is the presentation of REFLUX. In our fits (with both relxill and REFLUX), we have assumed that the spectrum illuminating the disk, and responsible for the reflection component can be described by a power law with a high-energy cutoff. However, this may not be the case for Swift J0243.6+6124. A power law with a high-energy cutoff is the spectrum expected for Comptonized photons, which is the case in which thermal photons from an accretion disk can inverse Compton scatter off free electrons in a hot corona. In the case of a neutron star, reflection features may be produced by illumination of the disk by thermal photons from the surface of the neutron star. Such a scenario could be implemented in REFLUX by using the xillverNS table, in which reflection spectra are produced by an incident blackbody spectrum.
While for neutron star XRBs like Swift J0243.6+6124 the Schwarzschild metric should describe well the spacetime geometry around the source, in the case of other super-Eddington X-ray sources (black hole XRBs, AGNs, and TDEs) deviations from the Schwarzschild background can be important and it will be necessary to implement the Kerr metric in REFLUX. This is also left to future work.
@@
This work was supported by the National Natural Science Foundation of China (NSFC), Grant No. 12250610185 and 12261131497, and the Natural Science Foundation of Shanghai, Grant No. 22ZR1403400. J.J. acknowledges support from Leverhulme Trust, Isaac Newton Trust and St Edmund’s College, University of Cambridge.

### v2

@@ 3 Super-Eddington TDE Swift J1644+57
TBabs $N_{H}$ $10^{22}$ cm-2 $^{*} 0.01730.0173^{*}$ $0.0173^{*}$
zTBabs $N_{H}$ $10^{22}$ cm-2 $1.557^{+0.049}_{-0.044}$ $1.597^{+0.055}_{-0.059}$ $1.443^{+0.029}_{-0.028}$
zcutoffpl $\Gamma$ $1.923^{+0.042}_{-0.051}$ $2.010^{+0.079}_{-0.087}$ $1.748^{+0.020}_{-0.020}$
$E_{\rm cut}$ keV $^{*} 100100^{*}$ $100^{*}$
norm $03440^{+0.0016}_{-0.0019} 0.0$$3680^{+0.0015}_{-0.0033} 0.0$$02874^{+0.0009}_{-0.0009}$
reflection $h$ $M$ $5.9^{+0.9}_{-1.5}$ $2.00^{+0.003}_{-P}$
$a$ $0.801^{+0.148}_{-0.410}$
$\theta_{\rm half}$ deg $29.133^{+5.754}_{-4.434}$
$v_{\infty}$ $c$ >0.384>0.384
$R_{\rm acc}$ $M$ $10^{*}$
$i$ deg $23.82^{P~\dagger}_{-8.71}$ $68.0^{+P}_{-1.8}$
$\log\xi$ erg cm s-1 <0.51<0.51 $0.58^{+0.50}_{-P}$
$A_{\rm Fe}$ solar <0.67<0.67 $0.50^{+0.08}_{-P}$
norm $\times 10^{-5}$ $1.294^{+6.503}_{-0.385}$ $376.3^{+1160.0}_{-28.5}$
Gaussian $E_{\rm line}$ keV $8.028^{+0.076}_{-0.078}$
$\sigma$ keV $0.0595^{+0.127}_{-P}$
norm $\times 10^{-5}$ $4.34^{+2.39}_{-2.02}$
$\chi^{2}/\nu$ $108.27/124$ $111.82/125$ $143.34/128$
Table 1: Redshift ($z$) in all models is set to $0.3534$. denotes the parameter was frozen and $P$ denotes the parameter reached its lower/upper limit. The galactic hydrogen column density ($N_{H}$) in model TBabs was obtained from https://www.swift.ac.uk/analysis/nhtot/index.php. The upper bound of inclination angle is always set to $\theta_{\rm half}$ in refluxlp_wind. The reported uncertainties correspond to the 90% interval limits for one relevant parameter ($\Delta\chi^{2}=2.71$).
@@ 4 Discussions
We tested our wind-reflection framework on the super-Eddington TDE Swift J1644+57. This source, powered by a tidal disruption of a star around a supermassive black hole, shows a highly super-Eddington accretion flow and has long been proposed as a case where powerful winds and relativistic jets coexist. The spectrum obtained with XMM-Newton presents clear reflection signatures in the Fe K band that cannot be adequately described by a simple Gaussian or a thin-disk reflection.
Our analysis with refluxlp_wind suggests a wind funnel with an half-opening angle of $\sim 30^{\circ}$. The wind is found to be very fast, with terminal velocity $v_{\infty}\gtrsim 0.384\,c$, and the system is observed at a low inclination $i\approx 24^{\circ}$. The corona is compact, located at $h\sim 6~M$, which allows for effective illumination of the funnel walls while still remaining within the gravitational well of the black hole. These geometric and kinematic constraints together explain why the observed Fe K feature is moderately blueshifted (as can also be seen in Fig. 5). The photons are scattered and reprocessed by the optically thick wind before escaping and the line-of-sight is within the funnel region. The opening angle and wind velocity might be related to the accretion rate as suggested from numerical simulations in Thomsen et al. (2019). We are not able to constrain the wind acceleration radius ($R_{\rm acc}$), possibly due to weak dependence of the iron line on the parameter (Fig. 5 top-right panel) and the data quality. Interestingly, we notice that the wind in non-ionized ($\log\xi\ \lesssim 0.51$) and has very low iron abundance ($A_{\rm Fe}\lesssim 0.67$). Just for comparison and testing, we also fit the data with relxilllp and a Gaussian.
Out of the three, the lowest $\Delta\chi^{2}$ is obtained with refluxlp_wind. The fitting results with relxilllp seem unphysical as the thin-disk geometry in the model predicts a high inclination of $i\sim 68^{\circ}$. At such a high inclination, the super-Eddington system should be obscured either by the slim-disk or the optically thick outflow. It also predicts extremely low corona height $h\sim 2M$ and a spin of $a\sim 0.8$. However, relxilllp fitting also predicts the low ionization and low iron abundance in the gas – consistent with the wind model. The Gaussian line fitting reveals only a narrow line ($\sigma\sim 0.06$ keV) located at $\sim 8$ keV with more residuals (Fig. 7). The fitting results of zTBabs and zcutoffpl are consistent across the three fitting results, with only the fit with zGauss having slight discrepancy (see Tab. 1).
From an astrophysical perspective, the wind geometry inferred for Swift J1644+57 is consistent with hydrodynamic and magnetohydrodynamic (MHD) simulations of super-Eddington accretion, which predict fast, optically thick outflows launched from the inner disk. The funnel angle and velocity we measure are comparable to those suggested by radiation-MHD models. A misclassification of wind reflection as disk reflection would bias spin or inclination estimates, highlighting the necessity of using models that capture the relevant geometry.
A more detailed analysis of the 2011 XMM-Newton data of Swift J1644+57 is beyond the purpose of the present work, which is the presentation of reflux. The wind-reflection fit of Swift J1644+57 emphasizes that super-Eddington systems can show fundamentally different reflection morphologies depending on whether the inner flow geometry is dominated by a geometrically thick disk or by a massive outflow. In all our fits, we have assumed that the spectrum illuminating the disk and responsible for the reflection component can be described by a power law with a high-energy cutoff. However, in the case of a TDE the reflection might be produced from the returning radiation of the thermal component from the disk, in such a case using xillverTDE table might be useful (Yao et al., 2024) but even better would be to have a reflection model which accounts for full returning radiation of thermal emissions (e.g. Mirzaev et al., 2024b, but with a super-Eddington geometry). Nevertheless, the TDE Swift J1644+57 was observed to have strong jet (Burrows et al., 2011), in such a case the base of the jet could act as the corona and our assumption stands well.
Our work motivates a broader application of both the slim-disk and wind-reflection models to a wide range of super-Eddington sources. These models applied with other spectral, timing and polarimetric data could really help decode the nature of accreting matter in extreme regimes. In future work, we plan to analyze data of sources which could possibly have the accretion geometry as in reflux. In terms of modeling, we wish to update reflux with the Kerr metric opening up the possibility of measuring black hole spins. We also hope to update the geometries with precise numerical accretion models.
@@
We would like to thank the anonymous reviewer for their constructive comments to improve the manuscript content. This work was supported by the National Natural Science Foundation of China (NSFC), Grant Nos. 12250610185, W2531002, and 12261131497. $S.$ acknowledges support from Shanghai Super Postdoctoral fellowship. $J.$ acknowledges support from Leverhulme Trust, Isaac Newton Trust and St Edmund’s College, University of Cambridge. Part of the computations in this research were performed using the CFFF platform of Fudan University.